OpenStudy (anonymous):

1 + 2 + 4 + 8 + 16 + … + 1,024?

OpenStudy (anonymous):

WELCOME TO OPEN STUDY! It would appear your missing some information. Is there more to the question?

OpenStudy (tkhunny):

Secret. Check it out. 1+2 = 3 = 4-1 1+2+4 = 7 = 8-1 1+2+4+8 = 15 = 16-1 1+2+4+8+16 = 31 = 32-1 1+2+4+8+16+32 = 63 = 64-1 Draw any conclusions from that?